OpenStudy (anonymous):

how do i write -4y=x-18 in slope intercept form?

OpenStudy (***[isuru]***):

slope intercept form y = mx + b where m is the slope and b is the intercept so... -4y = x - 18 4y = 18 - x y = (-1/4) x + 18/4 hope this will help ya!!